### v1

@@
XXII. Accurate stellar velocity dispersions of the SL2S lens sample and the lensing mass fundamental plane.
@@
(Received xxx, xxxx; accepted xxx, xxxx)
We reanalyze spectra taken as part of the SL2S lens galaxy survey, with the goal to obtain stellar velocity dispersion with precision and accuracy sufficient for time-delay cosmography. In order to achieve this goal, we impose stringent cuts on signal-to-noise ratio (SNR), and employ recently developed methods to mitigate and quantify residual systematic errors due to template libraries and fitting process. We also quantify the covariance across the sample. For galaxies with spectra with SNR >20/>20/> 20 /Å, our new measurements have average random uncertainty of 3-4%, average systematic uncertainty of 2%, and covariance across the sample of 1%. We find negligible covariance between spectra taken with different instruments. The systematic uncertainty and covariance need to be included when the sample is used as an external dataset in time-delay cosmography. We revisit empirical scaling relations of lens galaxies based on the improved kinematics. We show that the SLS2 sample, the TDCOSMO time-delay lens sample, and the lower redshift SLACS sample follow the same correlation between effective radius, stellar velocity dispersion and lensing mass, known as the lensing mass fundamental plane, as that derived by Auger et al. (2010) assuming isothermal mass profiles for the deflectors. We also derive for the first time the lensing mass fundamental plane assuming free power-law mass density profiles, and show that the three samples also follow the same correlation. This is consistent with a scenario in which massive galaxies evolve by growing their radii and mass, but staying within the plane.
Key Words.:
cosmology: distance scale – gravitational lensing: strong – Galaxy: kinematics and dynamics – Galaxies: elliptical and lenticular, cD
1 Introduction
The Hubble tension is arguably one of the most pressing open problems in cosmology (e.g., Di Valentino et al. 2025, and references therein.). The tension between the Hubble Constant H0 inferred from early and late universe probes has surpassed the canonical 5-$\sigma$ threshold for a statistically significant discrepancy. If the tension is not due to yet unidentified systematic errors, it implies that new physics is required, beyond the canonical flat $\Lambda$CDM.
@@ 1 Introduction
In order to quantify whether the tension is due to systematic errors, methods to measure H0 independent of the traditional local distance ladder and the sound horizon at recombination are especially valuable. Among these, the so-called time-delay cosmography stands out (Refsdal 1964; Treu et al. 2016, 2022; Birrer et al. 2024). It is completely independent of all other methods, it measures distances that are large enough to be unaffected by local over/underdensities and peculiar motions, and yet are small enough to be regarded as a late-time probe.
@@
State-of-the-art time-delay cosmography relies on stellar velocity dispersion measurements of the galaxies acting as deflectors (Treu & Koopmans 2002; Koopmans et al. 2003; Shajib et al. 2018; Yıldırım et al. 2020; Birrer et al. 2020; Shajib et al. 2023) to break the mass-sheet degeneracy (Falco et al. 1985). However, the requirements for stellar velocity dispersion measurements for time-delay cosmography are much more stringent than for galaxy evolution studies. The relative error on H0 is approximately double the relative error on stellar velocity dispersion, and thus one needs to control systematic errors and covariances between measurements at the percent level or better in order to achieve the 2% precision and accuracy required to distinguish between the Planck (Planck Collaboration et al. 2020) and SH0ES measurements (Riess et al. 2022). Motivated by this requirement, our team carried out a detailed investigation of stellar velocity dispersion kinematics. Knabel et al. (2025) showed that using carefully curated stellar template libraries and methods it is possible to achieve sub-percent accuracy, if spectra of sufficient SNR and resolution are available.
@@
In this paper, we re-analyze the spectra taken a decade ago as part of the SL2S survey (Ruff et al. 2011; Gavazzi et al. 2012; Sonnenfeld et al. 2013a, b, 2014, 2015) with the goal of implementing the recent improvements in the derivation of stellar kinematics, obtain accurate measurements and estimates of systematic errors and covariance between galaxies. SL2S is a very valuable external dataset for cosmography as it is well matched to the time-delay lenses in terms of redshift distribution (Birrer & Treu 2021). However, while 5-10% precision obtained by Sonnenfeld et al. (2013a) was sufficient to meet the main goal of the SL2S papers, i.e. to characterize the cosmic evolution of luminous and dark matter in early-type galaxies (Sheu et al. 2025, submitted), using them as external datasets for cosmography requires additional effort and more modern techniques.
@@
We also take advantage of the new measurements to investigate whether the SL2S sample galaxies obey the same lensing mass fundamental plane as the lower-redshift SLACS sample (Bolton et al. 2007, 2008b; Auger et al. 2010). The lensing mass plane is an empirical correlation similar to the traditional fundamental plane (Dressler et al. 1987; Djorgovski & Davis 1987), replacing surface brightness with total surface mass density obtained via the lens model. The lensing mass fundamental plane is tighter than the traditional fundamental plane and closer to the simple ”virial” relation, indicating a high degree of similarity between the mass distribution and kinematic structure of massive elliptical galaxies (Nipoti et al. 2008). We use the canonical lensing fundamental plane based on the assumption of isothermal mass density profile, and we introduce for the first time a new plane based on the lensing-only slope measurements to extrapolate the surface mass density to the standard circular aperture of radius equal to half the effective radius. We show that the SL2S and SLACS lenses follow the same planes as the time-delay lenses of the TDCOSMO sample (Millon et al. 2020; Birrer et al. 2020, and TDCOSMO collaboration 2025, in preparation). This finding is consistent with the assumption that external datasets have the same mass-structure as time-delay lenses necessary for obtaining joint constraints (Birrer et al. 2020; Birrer & Treu 2021).
@@
In Section 2, we summarize the SL2S spectroscopic data. In Section 3 we describe our methodology. In Section 4 we present our new measurements of stellar velocity dispersions and their uncertainties. In Section 5 we compare our new measurements with the previously published ones, and we study the lensing mass fundamental plane of the SL2S sample in comparison with SLACS and TDCOSMO. In Section 6, we summarize and discuss our results. When necessary, we adopt a flat $\Lambda$ cold dark matter cosmology, with $\Omega_{\rm m}=0.3$, $\Omega_{\Lambda}=0.7$ and $h=0.7$.
@@ 2 SL2S spectroscopic data
Spectroscopy for 58 SL2S candidates was obtained over several years with the goals of confirming the lensing hypothesis, securing source and deflector redshift, and measuring stellar velocity dispersion. Among these candidates, five were confirmed as non-lens, and the other 53 were analyzed by Sonnenfeld et al. (2013b). In this work, we excluded 6 of these 53 systems due to insufficient instrumental resolution and/or very low SNR, as described below in more detail. If multiple spectra were available for a galaxy from the same or different instrument, then we used the one with the highest SNR as our fiducial spectrum. We report also the results from the second-best spectra for completeness and we use them to perform a test of systematic errors. In this section, we briefly describe the three telescope/instrument setups that yielded the spectra of the galaxies in this sample. For specific details of the observations such as observing date, slit width, total exposure, and seeing, etc., check Table 2 in Sonnenfeld et al. (2013b).
@@ 2.1 Keck-LRIS
Among the 47 lens galaxies in this sample, 31 were observed with the Low-Resolution Imaging Spectrometer (LRIS; Oke et al. 1995) on Keck. LRIS consists of two detectors (blue and red) covering a wavelength range of 3000-10000 Å. The spectral resolution and reciprocal dispersion depend on the configuration of the grism (blue side) and grating (red side) used. The spectra were collected in the long-slit mode using both arms simultaneously with different exposure times. The observing nights were spread over several years, from 2006 to 2012, with typical seeing conditions (FWHM) ranging $\sim$ 0.6 ′′ - 1.3 ′′. In this paper, we only used spectra from the red side as the instrumental resolution for the setup used on the blue side was very low, prioritizing wavelength coverage for redshift determination. The 31 spectra with sufficient spectral resolution and SNR to determine stellar velocity dispersions are shown in Figures 1 and 2.
Figure 1: Spectra of the 16 galaxies from the 31 observed with LRIS on the Keck Telescope. In each panel, the black line is the observed spectrum, and the red line is the best-fit model from pPXF. The gray regions mark the excluded wavelength range from fitting and typically correspond to atmospheric absorption lines. Each panel also indicates the name of the system, redshift of the galaxy, and prominent absorption features present in the spectra.
Figure 2: As figure 1 for the remaining 15 of the 31 galaxies observed with LRIS.
@@ 2.2 VLT-Xshooter
The Xshooter spectrograph on VLT (Vernet et al. 2011) was used to obtain data for 13 systems. Xshooter consists of three independent cross-dispersed echelle spectrographs in three bands (UVB, VIS, NIR) and covers a huge wavelength range of 3000 - 25000 Å. This instrument has a higher resolution than LRIS and DEIMOS. The observations were performed using long slits of width 1.0′′ for the UVB arm, and 0.9′′ for the VIS and NIR arm. The observing nights were spread over three years from 2010 - 2012. The 13 spectra with sufficient spectral resolution and SNR to determine stellar velocity dispersions are shown in Figure 3.
Figure 3: Spectra of 13 galaxies observed with the Xshooter spectrograph on the VLT. In each panel, the gray line shows the observed spectrum, and the red line presents the best-fit model from pPXF. We have also plotted the smoothed version (the yellow line) of the observed spectrum. The gray bands mark the excluded wavelength range from fitting and typically correspond to the atmospheric absorption lines. Each panel also indicates the name of the system, redshift of the galaxy, and prominent absorption features present in the spectra.
2.3 Keck-DEIMOS
Three systems in this sample have data obtained with DEIMOS at Keck Observatory (Faber et al. 2003). The observations were performed using a long slit of width 1′′ with 600ZD grating. This setup covers a wavelength range of 4500 - 9500 Å with reciprocal dispersion of 0.65 Å. The typical instrumental resolution is around $\sigma_{\rm inst}\sim$ 70 km s-1.The spectra were collected during two observing nights in 2012. The 3 spectra with sufficient spectral resolution and SNR ratio to determine stellar velocity dispersions are shown in Figure 4.
Figure 4: Spectra of 3 galaxies from the SL2S sample observed with DEIMOS on the Keck Telescope. In each panel, the black line is the observed spectrum, and the red line is the best-fit model from pPXF. The gray regions mark the excluded wavelength range from fitting.
3 Methodology for new kinematic measurements
@@ 3.1 Summary of the method
We apply the recipe proposed by Knabel et al. (2025, hereafter KM25) to measure stellar kinematics and minimize and estimate the associated systematic uncertainties. First, we measured the instrumental line spread function (LSF), $\sigma_{\text{inst}}$ of the observed galaxy spectra. We also measured the SNR of each spectrum and, using simulations, determined the SNR required for each specific instrument to keep any bias below the 1% level. Using the information from KM25, we constructed two clean stellar libraries – Indo-US [^1] https://noirlab.edu/science/observing-noirlab/observing-kitt-peak/telescope-and-instrument-documentation/cflib (Valdes et al. 2004), and MILES [^2] https://miles.iac.es/ (Sánchez-Blázquez et al. 2006; Falcón-Barroso et al. 2011), customized for the wavelength range used in the kinematics fitting, as described in more detail below.
@@
Second, improper flux calibration, variance in continuum strength, etc., may bias the fit significantly if not corrected for using additive and multiplicative polynomials. To find optimal orders of the polynomials and strike a balance between under- and overfitting, we ran a grid of additive and multiplicative polynomials. We find that for low polynomial orders the fits are generally unstable, leave significant residuals, and result in disagreements between the template libraries, likely arising from improper flux calibration. Above a certain threshold, which depends on the wavelength range and on the quality of the flux calibration, the result becomes stable and yields good agreement between the templates. If the polynomial order is too high, the polynomials start modifying the shape of the absorption lines, and the fit becomes unstable again. So, it is important to choose polynomial orders in the ”sweet spot”, as discussed by Knabel et al. (2025). For this dataset, we find that orders of 8 and 2 for additive and multiplicative polynomials respectively, provide the least systematic scatter between templates, although similar results are found for somewhat lower and higher orders.
@@
Third, we select the wavelength range to fit. We intend to use the maximum possible wavelength range available for a galaxy spectrum to maximize the information content without compromising the SNR.
@@
Fourth, following the KM25 recipe, we use the Bayesian information criterion (BIC) to estimate the weighted average velocity dispersion ($\bar{\sigma}$) and associated systematic ($\Delta\bar{\sigma}$) and statistical uncertainties ($\delta\bar{\sigma}$). An important step in accomplishing this goal is to measure the threshold level for stellar templates to be considered as a significant contribution to the fit. If the threshold level is 1%, that means the templates with weight (w) $0.010.0<w<$ 0.01 are not contributing significantly to the fit (improving $\chi^{2}$) and should not be considered when counting the number of templates used to create the model. We measured the BIC of the fits for the whole sample, the BIC weight of each stellar library, and the BIC-weighted velocity dispersion and uncertainties using the equations provided by KM25.
@@
Figure 5: Example of typical variation in $\sigma_{\text{inst}}$ as a function of wavelength for three instruments – LRIS-red, Xshooter, and DEIMOS. In the plot for each instrument, there are two panels. The top one showing $\sigma_{\text{inst}}$ in km s-1 unit and the bottom one in Å unit over observed wavelength. In all panels the black line shows the average $\sigma_{\text{inst}}$ over the wavelength range.
3.2 Determination of instrumental resolution and threshold

### v2

@@
XXI. Accurate stellar velocity dispersions of the SL2S lens sample and the fundamental plane of the lensing mass
Pritom Mozumdar    Shawn Knabel These authors contributed equally and are considered co-first authors.    Tommaso Treu    Alessandro Sonnenfeld    Anowar J. Shajib    Michele Cappellari    Carlo Nipoti
(Received xxx, xxxx; accepted xxx, xxxx)
We reanalyzed spectra that were taken as part of the SL2S lens galaxy survey with the goal to obtain the stellar velocity dispersion with a precision and accuracy sufficient for time-delay cosmography. In order to achieve this goal, we imposed stringent cuts on the signal-to-noise ratio (S/N), and employed recently developed methods to mitigate and quantify residual systematic errors that are transferred from template libraries and fitting process. We also quantified the covariance across the sample. For galaxy spectra with S/N >20/>20/Å, our new measurements have an average random uncertainty of 3-4%, an average systematic uncertainty of 2%, and a covariance across the sample of 1%. We find a negligible covariance between spectra taken with different instruments. The systematic uncertainty and covariance need to be included when the sample is used as an external dataset in time-delay cosmography. We revisited empirical scaling relations of lens galaxies based on the improved kinematics. We show that the SL2S sample, the TDCOSMO time-delay lens sample, and the lower-redshift SLACS sample follow the same correlation of the effective radius, stellar velocity dispersion, and lensing mass, known as the lensing-mass fundamental plane, as the previously derived correlation that assumed isothermal mass profiles for the deflectors. We also derived for the first time the lensing-mass fundamental plane assuming free power-law mass density profiles, and we show that the three samples also follow the same correlation. This is consistent with a scenario in which massive galaxies evolve by growing their radii and mass, but stay within the plane.
Key Words.:
cosmology: distance scale – gravitational lensing: strong – Galaxy: kinematics and dynamics – Galaxies: elliptical and lenticular
1 Introduction
@@
In order to quantify whether the tension is caused by systematic errors, methods for measuring H0 independently of the traditional local distance ladder and the sound horizon at recombination are especially valuable. Among these, the so-called time-delay cosmography stands out (Refsdal64; Treu16; Treu22; Shajib24). It is completely independent of all other methods, it measures distances that are large enough to not be affected by local over or underdensities and peculiar motions, but are small enough to be regarded as a late-time probe.
@@
The most advanced time-delay cosmography relies on measurements of the stellar velocity dispersion of galaxies that act as deflectors (treu02b; koopmans03; shajib18; Yildirim20; birrer20; Shajib23) to break the mass-sheet degeneracy (Falco85). The requirements for measurements of the stellar velocity dispersion for time-delay cosmography are much more stringent than for galaxy evolution studies, however. The relative error on H0 is approximately twice the relative error on the stellar velocity dispersion, and the systematic errors and covariances between measurements therefore need to be controlled at the percent level or better in order to achieve the 2% precision and accuracy required to distinguish between the Planck (Planck_collab_2018) and ‘Supernovae, HO, for the Equation of State of Dark energy’ (SH0ES) measurements (Riess22). Motivated by this requirement, our team carried out a detailed investigation of the systematics of the stellar velocity dispersion in Knabel_Mozumdar_2025 and showed that a subpercent accuracy can be reached with curated stellar template libraries and methods when spectra with a sufficient S/N and resolution are available.
@@
In this paper, we reanalyze the spectra that were taken a decade ago as part of the ‘Strong Lensing Legacy Survey’ (SL2S) (Ruff11; Gavazzi12; Sonnenfeld_2013a; Sonnenfeld13b; Sonnenfeld14; Sonnenfeld15) with the goal of implementing the recent improvements in the derivation of stellar kinematics to obtain accurate measurements and estimates of the systematic errors and covariance between galaxies. SL2S is a very valuable external dataset for cosmography because it is well matched to the time-delay lenses in terms of redshift distribution and provides information that is helpful for breaking the mass-sheet degeneracy (B+T21), as demonstrated in a recent milestone paper by the TDCOSMO collaboration (Milestone25). While the precision of 5-10% obtained by Sonnenfeld13b was sufficient to meet the main goal of the SL2S papers, that is, to characterize the cosmic evolution of luminous and dark matter in early-type galaxies (William_2025_dinos_2), additional effort and more modern techniques are required to use them as external datasets for cosmography, however.
@@
We also took advantage of the new measurements to investigate whether the SL2S sample galaxies obey the same lensing-mass fundamental plane as the lower-redshift ‘Sloan Lens ACS Survey’ (SLACS) sample (Bolton07; Bolton08b; Auger_2010). The lensing-mass fundamental plane is an empirical correlation similar to the traditional fundamental plane (Dressler1987; DD1987). However, it replaces the surface brightness component of the traditional fundamental plane with the total surface mass density obtained via the lens model. The lensing-mass fundamental plane is tighter than the traditional fundamental plane and closer to the simple virial relation, which indicates a high degree of similarity between the mass distribution and kinematic structure of massive elliptical galaxies (NTB08). We used the canonical lensing-mass fundamental plane based on the assumption of an isothermal mass density profile, and we introduce for the first time a new plane that is based on the lensing-only slope measurements to extrapolate the surface mass density to the standard circular aperture with a radius equal to half the effective radius. We show that the SL2S and SLACS lenses follow the same planes as the time-delay lenses of the TDCOSMO sample (Milestone25; Millon_2020_TDCOSMO_I; birrer20). This finding is consistent with the assumption that external datasets have the same mass-structure as the time-delay lenses that are required to obtain joint constraints (birrer20; B+T21).
@@
In Sect. 2 we summarize the SL2S spectroscopic data. In Sect. 3 we describe our method. In Sect. 4 we present our new measurements of the stellar velocity dispersions and their uncertainties. In Sect. 5 we compare our new measurements with the previously published ones, and we study the lensing-mass fundamental plane of the SL2S sample in comparison with SLACS and TDCOSMO. In Sect. 6 we summarize and discuss our results. When necessary, we adopt a flat $\Lambda$ cold dark matter cosmology, with $\Omega_{\rm m}=0.3$, $\Omega_{\Lambda}=0.7$, and $h=0.7$.
@@ 2 SL2S spectroscopic data
The spectroscopy for 58 SL2S lens candidates was obtained over several years with the goals of confirming the lensing hypothesis, securing source and deflector redshift, and measuring the stellar velocity dispersion. Five of these candidates are not lens systems, and the other 53 were analyzed by Sonnenfeld13b. In this work, we excluded 6 of these 53 systems because the instrumental resolution was insufficient and/or the S/N was very low, as described below in more detail. When multiple spectra were available for a galaxy from the same or different instruments, we used the spectrum with the highest S/N as our fiducial spectrum. We also report the results from the second-best spectra for completeness, and we use them to perform a test of systematic errors.
In this section, we briefly describe the three telescope/instrument setups that yielded the spectra of the galaxies in this sample. For specific details of the observations, such as observing date, slit width, total exposure, and seeing, we refer to Table 2 in Sonnenfeld13b. The original HST images of the SL2S lens systems were presented by Sonnenfeld_2013a and were later updated with newly acquired observations by T24 and William_2025_dinos_2.
@@ 2.1 Keck-LRIS
Thirty-one of the 47 lens galaxies in this sample were observed with the Low-Resolution Imaging Spectrometer (LRIS; Oke95) on Keck. LRIS consists of two detectors (blue and red) that cover a wavelength range of 3000-10000 Å. The spectral resolution and reciprocal dispersion depend on the configuration of the grism (blue side) and grating (red side). The spectra were collected in the long-slit mode using both arms simultaneously with different exposure times. The observing nights were spread over several years, from 2006 to 2012, with typical seeing conditions (FWHM) ranging from $\sim$ 0.6 ′′ - 1.3 ′′. We only used spectra from the red side as the instrumental resolution for the setup used on the blue side was very low. The 31 spectra with a sufficient spectral resolution and S/N for determining the stellar velocity dispersions are shown in Figs. 1 and 2.
Figure 1: Spectra of the 16 of the 31 lens galaxies that were observed with LRIS on the Keck Telescope. In each panel, the black line shows the observed spectrum, and the red line shows the best-fit model from pPXF. The gray regions mark the wavelength range that we excluded from the fitting and typically these regions correspond to atmospheric absorption lines. Each panel also indicates the name of the system, the lens galaxy redshift, and prominent absorption features in the spectra.
Figure 2: Spectra of the remaining 15 of the 31 galaxies observed with LRIS. The line styles of the figure are same as Fig. 1.
2.2 Keck-DEIMOS
The data for three systems in this sample were obtained with the ‘DEep Imaging Multi-Object Spectrograph’ (DEIMOS; Faber_2003_DEIMOS) at Keck Observatory. The observations were performed using a long slit with a width of 1′′ with a 600ZD grating. This setup covers a wavelength range of 4500 - 9500 Å with a reciprocal dispersion of 0.65 Å. The typical instrumental resolution is about $\sigma_{\rm inst}\sim$ 70 km s-1. The spectra were collected during two observing nights in 2012. The three spectra with sufficient spectral resolution and S/N ratio for determining the stellar velocity dispersions are shown in Fig. 4.
Figure 3: Spectra of 13 lens galaxies observed with the Xshooter spectrograph on the VLT. In each panel, the gray line shows the observed spectrum, and the red line presents the best-fit model from pPXF. We also plot the smoothed version (the yellow line) of the observed spectrum. The gray bands mark the wavelength range we excluded from the fit and typically they correspond to the atmospheric absorption lines. Each panel also indicates the name of the system, the lens galaxy redshift, and prominent absorption features in the spectra.
@@ 2.3 VLT-Xshooter
The Xshooter spectrograph on VLT (Vernet_2011) was used to obtain data for 13 systems. Xshooter consists of three independent cross-dispersed echelle spectrographs in three bands (UVB, VIS, and NIR) and covers a huge wavelength range of 3000 - 25000 Å. This instrument has a higher resolution than LRIS and DEIMOS. The observations were performed using long slits with a width of 1.0′′ for the UVB arm, and 0.9′′ for the VIS and NIR arm. The observing nights were spread over three years from 2010 - 2012. The 13 spectra with a sufficient spectral resolution and S/N for determining the stellar velocity dispersions are shown in Fig. 3.
Figure 4: Spectra of three lens galaxies from the SL2S sample observed with DEIMOS on the Keck Telescope. In each panel, the black line shows the observed spectrum, and the red line shows the best-fit model from pPXF. The gray regions mark the wavelength range we excluded from the fit.
3 Method for new kinematic measurements
@@ 3.1 Summary of the method
We applied the recipe proposed by Knabel_Mozumdar_2025 to measure the stellar kinematics and minimize and estimate the associated systematic uncertainties. Using mock spectra with different S/N, KM25 assessed the recovery of the true velocity dispersion as well as the resulting uncertainty estimates and showed that the proposed method achieves the desired accuracy on mock samples. A detailed description of the simulation process is provided in the appendix of KM25. First, we measured the instrumental line spread function (LSF) $\sigma_{\text{inst}}$ of the observed galaxy spectra. We also measured the S/N of each spectrum and, using simulations, we determined the S/N required for each specific instrument to keep any bias below the 1% level. Using the information from KM25, we constructed two clean stellar libraries, Indo-US [^1] https://noirlab.edu/science/observing-noirlab/observing-kitt-peak/telescope-and-instrument-documentation/cflib (Valdes04), and MILES [^2] https://miles.iac.es/ (Sanchez_2006; Falcon-Barroso11), which were customized for the wavelength range we used in the kinematics fitting, as described in more detail below.
@@
Second, improper flux calibration, variance in continuum strength, and so on may bias the fit significantly when they are not corrected for using additive and multiplicative polynomials. To determine optimal orders of the polynomials and strike a balance between under- and overfitting, we ran a grid of additive and multiplicative polynomials. For low polynomial orders, the fits are generally unstable, leave significant residuals, and result in disagreements between the template libraries that likely arise from an incorrect flux calibration. Above a certain threshold that depends on the wavelength range and on the quality of the flux calibration, the result becomes stable and yields a good agreement between the templates. When the polynomial order is too high, the polynomials begin to modify the shape of the absorption lines, and the fit again becomes unstable. This shows that it is important to choose polynomial orders in the sweet spot, as discussed by KM25. For this dataset, orders of 8 and 2 for additive and multiplicative polynomials, respectively, provide the least systematic scatter between templates, although similar results were found for somewhat lower and higher orders.
@@
Third, we selected the wavelength range to fit. We intended to use the widest possible wavelength range available for a galaxy spectrum to maximize the information content without compromising the S/N.
Fourth, following the KM25 recipe, we used the Bayesian information criterion (BIC) to estimate the weighted average velocity dispersion ($\bar{\sigma}$) and the associated systematic ($\Delta\bar{\sigma}$) and statistical uncertainties ($\delta\bar{\sigma}$).
@@
An important step in accomplishing this goal was to measure the threshold level for stellar templates to be considered as a significant contribution to the fit. For a threshold level of 1%, the templates with a weight (w) $0.0<w<0.01$ do not contribute significantly to the fit (improving $\chi^{2}$) and should not be considered when the number of templates is counted that are used to create the model. We measured the BIC of the fits for the whole sample, the BIC weight of each stellar library, and the BIC-weighted velocity dispersion and uncertainties using the equations provided by KM25.
@@
Figure 5: Example of typical variation in $\sigma_{\text{inst}}$ as a function of wavelength for three instruments: LRIS-red, Xshooter, and DEIMOS. Each plot for each instrument has two panels. The top panel shows $\sigma_{\text{inst}}$ in km s-1 units, and the bottom panel shows it in Å unit over the observed wavelength. In all panels, the black line shows the average $\sigma_{\text{inst}}$ over the wavelength range.
3.2 Determining the instrumental resolution and threshold